OpenStudy (anonymous):

16^x=(1/512) how do you find x

OpenStudy (anonymous):

i would write both sides as a power of 2

OpenStudy (anonymous):

\[16=2^4\] so \[16^x=2^{4x}\]

OpenStudy (anonymous):

\[512=2^9\] so \[\frac{1}{512}=2^{-9}\]

OpenStudy (anonymous):

then you have \[2^{4x}=2^{-9}\] and therefore \[2x=-9\] etc

OpenStudy (anonymous):

you are greeaaat, im trying to review all of this for a quiz tomorrow, thank you!

OpenStudy (anonymous):

i can show you how to cheat with a calculator if you like. type is \[\log(\frac{1}{512})\div \log(16)\]

OpenStudy (anonymous):

myininaya where are you??